# Data-oriented Dynamic Fine-tuning Parameter Selection Strategy for FISH Mask based Efficient Fine-tuning

### v1

@@ Abstract
In view of the huge number of parameters of Large language models (LLMs) , tuning all parameters is very costly, and accordingly fine-tuning specific parameters is more sensible. Most of parameter efficient fine-tuning (PEFT) concentrate on parameter selection strategies, such as additive method, selective method and reparametrization-based method. However, there are few methods that consider the impact of data samples on parameter selecting, such as Fish Mask based method. Fish Mask randomly choose a part of data samples and treat them equally during parameter selection, which is unable to dynamically select optimal parameters for inconstant data distributions. In this work, we adopt a data-oriented perspective, then proposing an IRD (Iterative sample-parameter Range Decreasing) algorithm to search the best setting of sample-parameter pair for FISH Mask. In each iteration, by searching the set of samples and parameters with larger Fish information, IRD can find better sample-parameter pair in most scale. We demonstrate the effectiveness and rationality of proposed strategy by conducting experiments on GLUE benchmark. Experimental results show our strategy optimizes the parameter selection and achieves preferable performance. [^1] The code associated with this work will be made openly accessible following the anonymous review process.
Data-oriented Dynamic Fine-tuning Parameter Selection Strategy for FISH Mask based Efficient Fine-tuning
@@ 1 Introduction
LLMs have demonstrated outstanding capabilities in many fields, by performing auto-regressive training on massive Internet data. As a generalist model, LLMs are not specially optimized for one specific task during training. Therefore, supervised fine-tuning usually continues to improve the performance when facing specific problems. With the proposal of transfer learning, using a pre-trained model to fine-tune its parameters on the downstream task is becoming increasingly popular. However, as the parameter size of the pre-trained model keeps growing (e.g., GPT-3Brown et al. (2020), 175B parameters), it becomes challenging to fine-tune all parameters of LLMs. Consequently, various methods are proposed to alleviate the GPU memory consumption and training time by freezing most of the parameters in the neural network structure and only tuning some of them Lialin et al. (2023). This method named Parameter-Efficient Fine-Tuning (PEFT) or Delta TuningDing et al. (2023b).
@@
Several exemplary PEFT studies have been proposed, including but not limited to LoRA Hu et al. (2022), Adapter Houlsby et al. (2019), Prompt TuningLiu et al. (2022b), FISH Mask Sung et al. (2021), among others. These methodologies center on the design of new parameter structures or the subtly selection of specific parameters from existing networks. These existing works rarely pay attention to the impact of data on parameter selection before model training, even if it is necessary to use data for parameter selection. They just empirically select parameters or design additional structures, and then fine-tune the effectiveness of the method on the complete dataset.
Recently, there has been a focus on optimizing hyper-parameters in PEFT model from a data-oriented perspective, and the representative studies including Ding et al. (2023a); Valipour et al. (2023). This type of work focuses on how to optimize LoRA, a very representative method, which is difficult to extend to other structures. FISH Mask Sung et al. (2021) method is one of the few models that focuses on data-oriented non-LoRA structure optimization. It uses the Fisher-information matrix to determine the importance of parameters based on training data. However, this work assumes that all datasets are independent identically distributed, which is difficult to establish in real data.
To address this limitation, we propose IRD algorithm, which selects parameters refer to optimal samples rather than random samples. Through continuously reducing the scale of the set of parameters and sample, IRD can search the most effective sample-parameter pairs. The contributions of this work are summarized as follows:
  • We analyze FISH Mask method and find out that randomly selecting samples to calculate Fisher information Matrix is improper.
  • We propose an iterative range decreasing algorithm named IRD to optimize the representative PEFT method FISH Mask.
  • We conduct extensive experiments on a variety of tasks to demonstrate the effectiveness of our strategy.
@@
$\hat{F}_{\theta}$ $\hat{F}_{\theta}\in\mathbb{R}^{|\theta|}$ denotes the importance of each parameter.
@@
2 Related Work
2.1 Parameter Efficient Fine-tuning
With the increasing popularity of LLMs-related research, the PEFT method has developed rapidly in recent years. As introduced in Lialin et al. (2023), depending on the type of fine-tuning parameters, PEFT methods can be divided into three classes, including additive methods, selective methods, and reparametrization-based methods. Some representative additive methods include Adapter Houlsby et al. (2019), which involves adding small, fully-connected networks after attention and feed-forward network (FFN) layers in the existing Transformer models.
LoRA (Low-Rank Adaptation) Hu et al. (2022) is a reparametrization-based model. LoRA leverages the principle that neural networks can be effectively represented in low-dimensional spaces. This concept has received significant empirical and theoretical supportMaddox et al. (2020); Li et al. (2018); Arora et al. (2018); Malladi et al. (2023).
Selective methods do not add parameters or change the model structure, which fine-tune a small part of the parameters of the model. Diff pruning Guo et al. (2021) and BitFit Zaken et al. (2022) are two representative selective methods. Diff pruning learns a task-specific “diff” vector to extend the original parameters of the pretrained model. The diff vector is adaptively pruned during training to encourage sparsity. BitFit is a sparse fine-tune method, which only update the bias of the original model during tuning. BitFit can achieve competitive performance compared to dense full-parameter tuning.
2.2 Fisher Information
In deep learning neural networks, the concept of Fisher information has been extensively explored and applied. The Fisher information matrix (FIM) Fisher (1922); Amari (1996) plays a crucial role in mitigating the challenge of catastrophic forgettingFrench (1999); McCloskey and Cohen (1989); McClelland et al. (1995); Ratcliff (1990), especially in estimating the posterior distribution of task-specific parameters. The FIM possesses three key properties Pascanu and Bengio (2014) that make it highly effective in measuring the importance of parameters within a network. (a) FIM is analogous to the second derivative of the loss function at a minimum point. It is relevant to understand how tiny parameter perturbations can impact the network’s performance, especially around points of minimal loss. (b) FIM can be computed using first-order derivatives alone. This aspect simplifies the calculation process, making it feasible and less computationally intensive for models with a large number of parameters. (c) FIM is guaranteed to be positive semi-definite, ensuring that it describes a convex shape in the parameter space. This characteristic is important for maintaining stability in the optimization process and ensuring that parameter adjustments lead to predictable changes in the loss function. By quantifying the importance of parameters, the Elastic Weight Consolidation (EWC) can effectively balance learning new knowledge while preserving previously learned information.
@@ 3.2 Task Definition
The target of PEFT is to discern a subset of parameters for fine-tuning large models. The central task involves identifying or constructing a tunable parameter set. When the model, data, and final fine-tuning outcomes are identical, the fewer the parameters required for fine-tuning, the more effective is the method. In situations where the size of the fine-tuning parameter set is the same, the higher performance of the pre-trained model indicates the method is more effective.
3.3 FISH Mask based PEFT
FISH Mask is a sparse selective fine-tuning method introduced by Sung et al. (2021). This method employs Fisher information to select the top-k parameters of a model. FIM, defined as:
$\begin{split}&F_{\theta}=\mathbb{E}_{x\sim p(x)}[\\ &\mathbb{E}_{y\sim p_{\theta}(y|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\textrm{T}}]\end{split}$ (1)
where $x$ is the input, $y$ is the output, $\theta$ is the parameters of the model, $p(x)$ is the distribution of the input $x$. Fisher information is commonly estimated through a diagonal approximation, where gradients for all parameters are computed on several data batches:
$\hat{F}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{y\sim p_{\theta}(y|x_{i})}\left(\nabla_{\theta}\log p_{\theta}(y|x_{i})\right)^{2}$ (2)
In supervised learning, they use "empirical Fisher information" for further approximation:
$\hat{F}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\right)^{2}$ (3)
where $y_{i}$ is the ground-truth label for sample $x_{i}$. Once the top-k (empirical) Fisher information parameters are identified, only these parameters need to be optimized during fine-tuning. This process involves computing the (empirical) Fisher information for each parameter and creating a binary mask Zhao et al. (2020) based on a sparsity threshold, selectively updating only the most informative parameters in the mask.
@@ 3.3 FISH Mask based PEFT
FISH Mask also has some limitations. First, its initial phase is highly computationally intensive, necessitating the calculation of gradients for every parameter across numerous data batches. Second, it struggles with efficiency on modern deep learning hardware Lialin et al. (2023), which typically lacks support for sparse operations, leading to slower processing speeds. FISH Mask’s performance is inferior to methods like LoRA and $(IA)^{3}$ Liu et al. (2022a), even though it is on par with Adapters.
@@ 4.1 Intuition
For most of the selective-based PEFT methods, the performance of fine-tuning is related to the scale of the tuned parameters and the size of dataset, which is similar to the scaling-law theory of pretrain language model Kaplan et al. (2020). For FISH Mask, the fine-tuning performance is affected by the parameter scale(Mask Sparsity) and the number of tuning samples, as shown in Fig.1. Once the Mask Sparsity and the set of samples are fixed, FISH mask will select the optimal parameters according to the FIM of the set of samples. However, FISH Mask only focuses on the size of the sample set, and each element in the set is randomly selected. We think that this setting has drawbacks and does not consider the diversity of the samples. When the size of the sample set is determined, selecting better samples to form the set will effectively improve the effect of fine-tuning.
@@ 4.2 Data Driven Assumption
$\Theta_{0}^{*}=\{\theta_{i}\in\theta|\hat{F}_{\theta_{i}}>={\rm sort}(\hat{F}_{\theta})_{k}\}$ (4)
The setting of Eq.4 is following Sung et al. (2021), they assume that each training example $x_{i}$ in $\hat{F}_{\theta}$ is independent and identically distributed. In practical datasets, maintaining a consistent data distribution proves challenging, with the presence of considerable noise data making it further complicated. Therefore, there are limitations to random select samples to calculate FIM. In this work, we assume that the samples with larger Fisher-Information (See in Section2.2) are more important in fine-tuning, which will directly affect the quality of parameter selection.
4.3 Iterative Range Decreasing
We propose IRD algorithm to select better samples instead of random selecting. The inputs of IRD algorithm are model $\mathcal{M}$ to be fine-tuned, the set of initial input samples $X_{0}$, and the ground-truth $Y^{*}$. We firstly set variable $l$ to record the iteration and $S$ to receive the output (Step 1). With these inputs, IRD firstly generates the $\Theta_{0}^{*}$, which represents the initial set of parameters with top-k Fisher information (Step 2). Then, the iterative range decreasing process is conducted. During each iteration, if the number of elements in both $X_{l}$ and $\Theta_{l}^{*}$ is greater than one, we alternately halve the range of both $X_{l}$ and $\Theta_{l}^{*}$ (Step 4 and Step 7). And the performance of $\mathcal{M}$ of specific pair of sample-parameter is calculated once the sample or parameter set changes (Step 5 and Step 8). Finally, we get a set of performance scores (calculated by evaluation metrics in Section 5.1) and a corresponding set of tuned parameters.
Input: $X_{0}$; $Y^{*}$$\mathcal{M}$
Output: $S$
1 $l=0$; $S=\emptyset$; $\Theta_{0}^{*}=\{\theta_{i}\in\theta|\hat{F}_{\theta_{i}}>={\rm sort}(\hat{F}_{\theta})_{k}\}$; for $|X_{l}|>1$ and $|\Theta_{l}^{*}|>1$ do
2       $X_{l+1}=\{x_{i}\in X_{l}|\hat{F}_{x_{i}}>={\rm sort}(\hat{F}_{x_{i}})_{\frac{|X_{l}|}{2}}\}$; $s={\rm Score}(X_{l+1},Y^{*},\mathcal{M}_{\Theta_{l}^{*}})$; $S=S+\{(s,\Theta_{l}^{*})\}$; $\Theta_{l+1}^{*}=\{\theta_{i}\in\Theta_{l}|\hat{F}_{\theta_{i}}>={\rm sort}(\hat{F}_{\theta})_{\frac{|\Theta_{l}|}{2}}\}$ $s={\rm Score}(X_{l+1},Y^{*},\mathcal{M}_{\Theta_{l+1}^{*}})$; $S=S+\{(s,\Theta_{l+1}^{*})\}$; $l=l+1$

### v2

# 1]organization=Hubei Provincial Key Laboratory of Artificial Intelligence and Smart Learning, addressline=, city=Wuhan, postcode=430079, state=Hubei, country=PR China

@@ Abstract
Fine-tuning all parameters of Large Language Models (LLMs) is computationally expensive. Parameter-Efficient Fine-Tuning (PEFT) methods address this by selectively fine-tuning specific parameters. Most of the parameter efficient fine-tuning (PEFT) methods center on selecting or introducing a set of parameters to be fine-tuned. However, there are few methods that consider the impact of data samples on parameter selecting. Representative data driven methods include FISH Mask based method, which randomly selects a portion of data samples as a basis when selecting parameters. However, this random data sample selection method cannot select optimal parameters for unstable data distribution. In this work, we introduce a data-centric approach and propose the Iterative Range Decreasing (IRD) algorithm to optimize the sample-parameter pair selection in FISH Mask. IRD iteratively refines the selection by identifying subsets of samples and parameters exhibiting higher Fisher information. We demonstrate the effectiveness and rationality of proposed strategy by conducting experiments on GLUE benchmark. Experimental results show our strategy optimizes the parameter selection and achieves preferable performance over some typical baseline methods.
keywords:
@@ 1 Introduction
Large language models have demonstrated fabulous capabilities across various fields through auto-regressive training on vast datasets from the internet. As generalist models, they are not optimized for any specific task during training. Therefore, supervised fine-tuning usually continues to improve the performance when facing specific problems. With the proposal of transfer learning, using a pre-trained model to fine-tune its parameters on the downstream task is becoming increasingly popular. However, as the parameter size of the pre-trained model keeps growing (e.g., GPT [3], 175B parameters), it becomes challenging to fine-tune all parameters of LLMs. Consequently, various methods are proposed to alleviate the GPU memory consumption and training time by freezing most of the parameters in the neural network structure and only tuning some of them  [19]. This method named Parameter-Efficient Fine-Tuning (PEFT) or Delta Tuning [8].
@@
Several exemplary PEFT studies have been proposed, including but not limited to LoRA [14], Adapter [13], Prompt Tuning[20], among others. These methodologies center on the design of new additional parameter structures or the subtly selection of specific parameters from existing networks. These existing works rarely pay attention to the impact of data on parameter selection before model training, even if it is necessary to use data for parameter selection. They just empirically select parameters or design additional structures, and then fine-tune the effectiveness of the method on the complete dataset.
Data-oriented approaches [7, 32, 37] offer a promising direction for optimizing parameter selection in PEFT. One such method, FISH Mask [31], leverages the Fisher Information Matrix (FIM) to estimate the importance of each parameter with respect to the training data. Specifically, FISH Mask calculates the FIM based on a randomly selected subset of training data, using the FIM’s diagonal entries to rank parameter importance. This ranking then guides the selection of parameters to be fine-tuned. However, the assumption of independent and identically distributed data and the use of random sampling for FIM calculation in FISH Mask limit its effectiveness in real-world scenarios where data distributions can be complex and non-uniform. This raises a critical research question: How can we improve the sample selection process in FISH Mask to achieve better parameter selection and fine-tuning performance?
To address this limitation, we propose the Iterative Range Decreasing (IRD) algorithm, which iteratively refines the selection of samples and parameters based on their Fisher information. IRD identifies a sample set with higher Fisher information content, leading to more informed parameter selection. In summary, the contributions of this work are:
  • We investigated Parameter-Efficient Fine-Tuning (PEFT) methods based on selective strategies and observed that high-quality training samples are crucial for selecting optimal fine-tuning parameters. Analysis of the FISH Mask method revealed that random sample selection for calculating the Fisher Information Matrix (FIM) limits its performance.
  • To address this limitation, we propose an Iterative Range Decreasing (IRD) algorithm to optimize the sample selection process. IRD identifies a sample set with higher Fisher information, which we utilize for parameter selection during fine-tuning.
  • Extensive experiments on various representative benchmarks demonstrate the superiority of our proposed strategy. Using only 0.2% of model parameters for fine-tuning, our method significantly outperforms existing approaches, including widely used methods such as LoRA.
2 Related Work
2.1 Parameter Efficient Fine-tuning
With the increasing popularity of LLMs-related research, the PEFT method has developed rapidly in recent years. As introduced in Lialin et al. [19], depending on the type of fine-tuning parameters, PEFT methods can be divided into three classes, including additive methods, selective methods, and reparametrization-based methods. Some representative additive methods include Adapter  [13], which involves adding small, fully-connected networks after attention and feed-forward network (FFN) layers in the existing Transformer models.
Adapters have demonstrated their effectiveness across various NLP tasks, including the GLUE benchmark, where they nearly matched the performance of fully fine-tuned BERT models. LoRA (Low-Rank Adaptation) [14] is a reparametrization-based model, which leverages the principle that neural networks can be effectively represented in low-dimensional spaces. This concept has received significant empirical and theoretical support [21, 18, 2, 22]. LoRA utilizes a simple low-rank matrix decomposition to parameterize weight updates in neural networks, ensuring that pre-trained model weights remain frozen while only the newly added low-rank matrices are trainable.
Selective methods do not add parameters or change the model structure, which fine-tune a small part of the parameters of the model. Diff pruning  [12] and BitFit [38] are two representative selective methods. Diff pruning learns a task-specific “diff” vector to extend the original parameters of the pretrained model. The diff vector is adaptively pruned during training to encourage sparsity. BitFit is a sparse fine-tune method, which only update the bias of the original model during tuning. BitFit can achieve competitive performance compared to dense full-parameter tuning.
2.2 Fisher Information
In deep learning neural networks, the concept of Fisher information has been extensively explored and applied. The Fisher information matrix (FIM)  [10, 1] plays a crucial role in quantifying parameter importance, which is then leveraged by methods addressing the challenge of catastrophic forgetting  [11, 25, 24, 29]. While not directly mitigating forgetting, the FIM provides essential information about parameter sensitivity. The FIM possesses three key properties  [26] that make it highly effective in measuring the importance of parameters within a network: (a) Relationship to the Hessian: The FIM approximates the Hessian matrix of the loss function around a minimum point. This is crucial for understanding how small changes in parameters affect the network’s performance, especially near optimal points. It provides insights into the curvature of the loss landscape. (b) Computational Efficiency: The FIM can be computed using only first-order derivatives. This significantly simplifies the calculation process, making it feasible and less computationally intensive for LLMs with numerous parameters. (c) Positive Semi-Definiteness: The FIM is guaranteed to be positive semi-definite. This ensures it describes a convex shape in the parameter space, which is important for stability during optimization. This property ensures that parameter updates lead to predictable changes in the loss function. By quantifying the importance of parameters, methods like Elastic Weight Consolidation (EWC)  [26] can effectively balance learning new knowledge while preserving previously learned information. EWC leverages the FIM to identify parameters crucial for earlier tasks and applies stronger regularization to these parameters during subsequent training, preventing drastic changes and thus mitigating catastrophic forgetting.
@@ 2.2 Fisher Information
^$\hat{\mathcal{I}}_{\theta}\in\mathbb{R}^{|\theta|}$ Empirical FIM that denotes the importance of each parameter.
@@ 3.2 Task Definition
The goal of PEFT is to select a subset of parameters for fine-tuning LLMs, rather than fine-tuning all parameters. This involves identifying a subset of existing parameters, or constructing a new set of tunable parameters (e.g., as in LoRA). Two key factors determine the effectiveness of a PEFT method: the number of tunable parameters and the final performance. Ideally, a PEFT method achieves comparable performance to full fine-tuning with significantly fewer parameters. Therefore, given the same model, data, and fine-tuning objective, a smaller parameter set leading to higher performance indicates a more effective method.
3.3 Fisher Information Matrix and FISH Mask based PEFT
FISH Mask is a sparse selective fine-tuning method introduced by  Sung et al. [31]. This method employs Fisher information to select the top-k parameters of a model. FIM is defined as:
$\mathcal{I}_{\theta}=\mathbb{E}_{x\sim p(x)}[\mathbb{E}_{y\sim p_{\theta}(y|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\textrm{T}}],$ (1)
where $x$ is the input, $y$ is the output, $\theta$ is the parameters of the model, $p(x)$ is the distribution of the input $x$. $\nabla_{\theta}$ represents the gradient of the objective function with respect to the model parameters $\theta$, computed based on the given data. Fisher information is commonly estimated through a diagonal approximation, where gradients for all parameters are computed on several data batches:
^$\hat{\mathcal{I}}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{y\sim p_{\theta}(y|x_{i})}\left[\nabla_{\theta}\log p_{\theta}(y|x_{i})\odot\nabla_{\theta}\log p_{\theta}(y|x_{i})\right],$ (2)
where $N$ is the number of data batches. $\odot$ is Hadamard old. In supervised learning, they use “Empirical Fisher information” for further approximation:
^$\hat{\mathcal{I}}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\odot\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})$ (3)
where $y_{i}$ is the ground-truth label for sample $x_{i}$. Once the top-k (empirical) Fisher information parameters are identified, only these parameters need to be optimized during fine-tuning. This process involves computing the (empirical) Fisher information for each parameter and creating a binary mask [39] based on a sparsity threshold, selectively updating only the most informative parameters in the mask.
@@ 4.1 Intuition
For most of the selective-based PEFT methods, the performance of fine-tuning is related to the scale of the tuned parameters and the size of dataset, which is similar to the scaling-law theory of pretrain language model [15]. For FISH Mask, the fine-tuning performance is affected by the parameter scale (Mask Sparsity) and the number of tuning samples, as shown in Fig.1. Once the Mask Sparsity and the set of samples are fixed, FISH mask will select the optimal parameters according to the FIM of the set of samples. However, FISH Mask only focuses on the size of the sample set, and each element in the set is randomly selected. We think that this setting has drawbacks and does not consider the diversity of the samples. When the size of the sample set is determined, selecting better samples to form the set will effectively improve the effect of fine-tuning.
@@ 4.2 Data Driven Assumption
$\Theta_{0}^{*}=\{\theta_{i}\in\theta|\hat{\mathcal{I}}_{\theta_{i}}>={\rm sort}(\hat{\mathcal{I}}_{\theta})_{k}\}$ (4)
The setting of Eq.4 is following Sung et al. [31], they assume that each training example $x_{i}$ in ^$\hat{\mathcal{I}}_{\theta}$ is independent and identically distributed. In practical datasets, maintaining a consistent data distribution proves challenging, with the presence of considerable noise data making it further complicated. Therefore, there are limitations to random select samples to calculate FIM. In this work, we assume that the samples with larger Fisher-Information are more important in fine-tuning, which will directly affect the quality of parameter selection.
Input: The set of initial chosen samples $X_{0}$; The set of initial chosen parameters $\Theta_{0}$; Ground truth of fintuned data $Y^{*}$; LLM to be tuned $\mathcal{M}$, $k$ is the number of initail parameters to be tuned.
Output: $S$: Optimal parameter sets and its corresponding performance after fine-tuning.
1 $l=0$; $S=\emptyset$;
2^$\hat{\mathcal{I}}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\odot\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})$
[…]
4for $|X_{l}|>1$ and $|\Theta_{l}^{*}|>1$ do
5       $X_{l+1}=\{x_{i}\in X_{l}|\hat{\mathcal{I}}_{x_{i}}>={\rm sort}(\hat{\mathcal{I}}_{x_{i}})_{\frac{|X_{l}|}{2}}\}$;
6      $s={\rm Score}(X_{l+1},Y^{*},\mathcal{M}_{\Theta_{l}^{*}})$;
7      $S=S+\{(s,\Theta_{l}^{*})\}$;
8      $\Theta_{l+1}^{*}=\{\theta_{i}\in\Theta_{l}|\hat{\mathcal{I}}_{\theta_{i}}>={\rm sort}(\hat{\mathcal{I}}_{\theta})_{\frac{|\Theta_{l}|}{2}}\}$
9      $s={\rm Score}(X_{l+1},Y^{*},\mathcal{M}_{\Theta_{l+1}^{*}})$;
10      $S=S+\{(s,\Theta_{l+1}^{*})\}$;
11      $l=l+1$
@@
4.3 Iterative Range Decreasing
We propose IRD algorithm to select better samples instead of random selecting. IRD optimizes first phase of FISH mask. The goal of IRD algorithm is to find the optimal set of parameters for fine-tuning. In each iteration, IRD reduces the scales of the set of data samples and the set of candidate parameters. The optimal parameter set obtained in each iteration is recorded. When IRD ends, we can obtain the optimal parameter set corresponding to different data scales from large to small.
As shown in Algorithm1, $l$ is the variable to control iteration. At the $l-$th iteration, $X_{l}$ is the set of data samples at and $\Theta_{l}^{*}$ is the set of parameters. $s$ is the fine-tuning performance based on the parameter set of the current iteration. The inputs of IRD algorithm are model $\mathcal{M}$ to be fine-tuned, the set of initial input samples $X_{0}$, and the ground-truth $Y^{*}$. We firstly set variable $l$ to record the iteration and $S$ to receive the output (Step 1). With these inputs, IRD firstly generates the $\Theta_{0}^{*}$, which represents the initial set of parameters with top-k Fisher information (Step 2). Then, the iterative range decreasing process is conducted. During each iteration, if the number of elements in both $X_{l}$ and $\Theta_{l}^{*}$ is greater than one, we alternately halve the range of both $X_{l}$ and $\Theta_{l}^{*}$ (Step 4 and Step 7). And the performance of $\mathcal{M}$ of specific pair of sample-parameter is calculated once the sample or parameter set changes (Step 5 and Step 8). Finally, we get a set of performance scores (calculated by evaluation metrics in Section 5.1) and a corresponding set of tuned parameters.
4.4 Theoretical Insights into Sample Selection
While a rigorous mathematical proof of the superiority of our proposed sample selection method is complex and depends on the specific model architecture and data distribution, we offer a simplified example to illustrate why non-random sample selection strategies can be advantageous.
4.4.1 Simplified Linear Model
Consider a linear model $y=\theta^{T}x+\epsilon$, where $x\in\mathbb{R}^{d}$ is the input, $y$ is the output, $\theta\in\mathbb{R}^{d}$ is the parameter vector, and $\epsilon\sim N(0,\sigma^{2})$ is Gaussian noise. Assuming $x\sim N(0,I)$, the Fisher information matrix simplifies to $\mathcal{I}(\theta)=\frac{1}{\sigma^{2}}I$. Our goal is to select parameters that have the largest impact on the output prediction, which can be interpreted as selecting parameters along the eigenvector corresponding to the largest eigenvalue of the Fisher information matrix.
4.4.2 Random Sample Selection
With $n$ randomly selected samples $\{x_{1},x_{2},...,x_{n}\}$, the empirical Fisher information matrix is ^$\hat{\mathcal{I}}(\theta)=\frac{1}{n\sigma^{2}}\sum_{i=1}^{n}x_{i}x_{i}^{T}$. Due to the randomness, ^$\hat{\mathcal{I}}(\theta)$ fluctuates around $\frac{1}{\sigma^{2}}I$, especially for small $n$. This large variance hinders accurate estimation of the true Fisher information, making it difficult to select important parameters effectively. Due to the randomness, ^$\hat{\mathcal{I}}(\theta)$ fluctuates around $\frac{1}{\sigma^{2}}I$. This is a consequence of sampling error: the empirical Fisher information is computed from a finite sample, while the true Fisher information is based on the entire population distribution. Specifically, ^$\hat{\mathcal{I}}(\theta)$ is a random matrix, as it is the average of the random matrices $x_{i}x_{i}^{T}$. By the law of large numbers, ^$\hat{\mathcal{I}}(\theta)$ converges to $\mathcal{I}(\theta)$ as $n\to\infty$. However, for finite $n$, ^$\hat{\mathcal{I}}(\theta)$ is merely an estimator and exhibits variability around $\mathcal{I}(\theta)$. This variability is particularly pronounced for small $n$, as a small sample set is less representative of the population and more susceptible to random fluctuations. This increased variance in ^$\hat{\mathcal{I}}(\theta)$ hinders accurate estimation of the Fisher information and thus impedes effective selection of important parameters.
4.4.3 Non-Random Sample Selection
Suppose we strategically select samples aligned with a specific direction $v$, such that $x_{i}\approx\alpha_{i}v$, where $\alpha_{i}$ is a scalar. The empirical Fisher information becomes ^$\hat{\mathcal{I}}(\theta)\approx\frac{1}{n\sigma^{2}}\sum_{i=1}^{n}\alpha_{i}^{2}vv^{T}$. The dominant eigenvector of this matrix aligns with $v$, indicating that we can amplify the Fisher information along $v$ by choosing samples aligned with it. This targeted selection facilitates more efficient selection of parameters related to this direction.
4.4.4 Advantages of Non-Random Sampling Strategies for Parameter Selection
Random sampling yields an unbiased estimate of $\mathcal{I}(\theta)$, but with potentially high variance, especially for small $n$. Non-random sampling such as IRD, while introducing bias, reduces the variance of the estimator. While random sampling offers an unbiased estimate of the Fisher Information matrix, $\mathcal{I}(\theta)$, non-random, or strategic, sampling presents distinct advantages for parameter selection, particularly when the objective is focused on a specific direction in parameter space. Non-random sampling allows for the targeted acquisition of information relevant to a predefined direction, $v$, in parameter space. By concentrating sampling efforts along $v$, the empirical Fisher information, ^$\hat{\mathcal{I}}(\theta)$, becomes more informative about the parameters that influence variations along this direction. This focused approach is particularly beneficial when prior knowledge or task-specific requirements dictate a specific direction of interest. For example, if we know that certain parameters are more likely to impact a specific performance metric, we can tailor our sampling strategy to prioritize those parameters. Non-random sampling can significantly reduce the variance of the estimator, ^$\hat{\mathcal{I}}(\theta)$, compared to random sampling, especially when the sample size is limited. This reduced variance enhances the sensitivity of the estimation to changes in the parameters along the chosen direction $v$, making it easier to identify the most influential parameters. This increased sensitivity allows for more confident selection of important parameters, even with fewer samples, potentially saving computational resources.
By prioritizing samples with high Fisher Information for subsequent iterations, IRD offers several key advantages. This strategy focuses on the most informative samples, those that contribute most to accurately estimating the Fisher Information matrix, $\mathcal{I}(\theta)$, leading to faster convergence and more precise parameter estimates with fewer samples. Furthermore, utilizing highly informative samples enhances the sensitivity of $\mathcal{I}(\theta)$ estimation, allowing for more accurate identification of influential parameters. Finally, this focused approach enables efficient resource allocation by concentrating computational effort on the most informative data points, resulting in significant computational savings, especially beneficial for large datasets or complex models.
4.5 Time and Space Complexity Analysis
IRD is a deep learning algorithm, and its time and space complexity is difficult to accurately estimate. The time complexity of IRD depends on the number of data samples $n=|X_{0}|$ and the parameter scale $m=|\Theta_{0}^{*}|$ of the foundation model $\mathcal{M}$. The rough time complexity is $\log_{2}nm$, where the dominant term is a fine-tuning process. The space complexity is $km$, which depends on the different optimizer. After the model and fine-tuning data set are determined, executing the IRD algorithm once can obtain multiple optimal fine-tuning parameter sets of different data scales.